• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 14:56:28 on March 8,2002

    Entry number 1195

    keyword=cryotarget cool-down completed

    Mike Seely started cooldown early today. He started with both loops (loop 1 for LH2
    and loop3 for LD2). Loop 1 is going to be used for the Keon experiment and Loop 3
    was to test if the fan works or not. Due to problem with cooling power, loop 3
    cooldown stopped before it could be completed. I took over the cool-down at 1 pm
    when the LH2 loop was at about 23K and it was in the middle of switching from
    warm return to cold return. Completed the switch at 1:10 pm. ESR supply temperature increased to >21K before it eventally came down to < 14K.
    Fully closed pre-cool J-T valve at 1:30 pm. Added some Hydrogen gas to the loop. and increased fan speed slowly to 60 Hz. By 2:30 pm the target cool-down completed.